Based on a​ poll, 60​% of adults believe in reincarnation. Assume that 6 adults are randomly​ selected, and find the indicated probability. What is the probability that all the selected adults believe in reincarnation


1
Expert's answer
2021-04-13T09:40:58-0400

"P(x=6)=C^6_6\cdot0.6^6\cdot(1-0.6)^0=0.0467"